OpenStudy (anonymous):

Translate into an equation and solve. The sum of a number, n, and eight is twelve less than six times the number.

OpenStudy (radar):

n+8 =6n-12 5n=20 n=4

OpenStudy (radar):

As you can see 4+8=24-12 12=12